Suppose you invest $6,000 and would like your investment to grow to $8,000 in five years. What interest rate, compounded monthly, would you have to earn in order for this to happen?
Answer by jorel1380(3719)

8000/6000=1.333
In order for your money to grow by 1/3 in 5 years, you need:
4/3=(1+r/12)^60
ln 4/3= ln(1+r/12)^60
ln 4/3=60 ln(1+r/12)
0.00479470120752968212398698343323=ln (1+r/12)
e^0.00479470120752968212398698343323 = e^ln (1+r/12)
1.0048062141804317245207735522828 = l+r/12
r=0.05767457016518069424928262739328
Your yearly interest rate needs to be about 5.7674% in order for your money to grow by 1/3 in 5 years.